# Detecting Visual Information Manipulation Attacks in Augmented Reality: A Multimodal Semantic Reasoning Approach

### v1

@@ 4.2 Data Collection Pipeline
Monitor-Based AR Data Collection: We started by collecting 58 high-resolution real-world scene images from online sources and generative AI models. The AI-generated backgrounds is a double-edged sword: they enable a broader range of scenarios that are often difficult to capture in the real world, but may also introduce domain shifts that may affect the behavior of VLMs. All of these backgrounds form a diverse set of environments such as street signs, building facades, hospital entrances, and parking instructions. We display these images on a monitor one by one as background. For each background, we manually designed virtual content to simulate either attack or non-attack cases. The content varies across attack formats and purposes, based on the taxonomy introduced in Section˜3. These assets were imported into Unity and positioned at designated regions. Then, we implemented an AR application for Android phones using Unity with ARCore support. Specifically, we utilized ARCore’s image tracking capability to precisely align the virtual content with the background scene. In our setup, each background image is used as a reference image. When the phone’s camera detects the reference image at runtime, ARCore registers its spatial pose and enables Unity to attach virtual content to the same coordinate system. After that, we recorded each AR experience using two virtual AR cameras in Unity. One camera renders only the background scene to produce the raw image , while the other renders both the background and the virtual content to produce the AR image . Finally, we combined sequences of raw/AR images into raw/AR video pairs, each pair representing a specific scenario.
@@
Real-World AR Data Collection: While the monitor-based pipeline allows for efficient prototyping and introduces a more diverse range of scenarios, it may fail to capture important spatial properties of real-world AR environments such as 3D depth cues, which are intrinsic to real AR environments. To address this limitation, we additionally collected some data in real-world settings. Specifically, we developed an AR application for the Meta Quest 3 in Unity 2022.3.61f1, which supports access to the device’s main camera. The app allows users to grab and place virtual content at desired locations using the controller. We manually designed 35 combinations of real-world scenes and virtual content to comprehensively cover all attack formats and purposes described in Section˜3. The main camera of the Meta Quest 3 was accessed to capture raw images , while a virtual camera in Unity rendered the virtual content separately. The virtual content was then overlaid onto the raw images to produce the corresponding AR images . Finally, we also combined image sequences to create the videos.
@@ 4.3 Dataset Composition
With the data collection pipelines, we developed the AR-VIM dataset. The final AR-VIM dataset consists of 452 video pairs, each containing a raw video and its corresponding augmented version. These pairs span a total of 202 unique scenes of AR experiences. Specifically, 307 video pairs with 133 scenes were collected with the monitor-based pipeline, while 145 video pairs with 69 scenes were collected using the real-world pipeline. Each pair simulates either an attack or non-attack scenario using the taxonomy introduced in Section˜3. For each video pair, the virtual content appears at a specific timestamp, simulating a real-time AR interaction where the virtual overlay is introduced during scene observation. Each video in the dataset has a resolution of either 480×1080 pixels (monitor-based videos) or 960×1280 pixels (real-world videos), and a duration ranging from 4 to 17 seconds. The videos had a framerate of 15 FPS, a constraint imposed by smartphone hardware limitations, as running an AR application while simultaneously capturing high-frame-rate raw and AR videos places strain on AR devices.
To ensure balanced coverage across both manipulation formats and purposes, we constructed seven valid types of attacks: character replacement, phrase replacement, phrase obfuscation, phrase extra information, pattern replacement, pattern obfuscation, and pattern extra information. For each attack type, we also generated a comparable number of non-attack cases for comparison. In total, the dataset includes 241 attacked video pairs and 211 non-attack video pairs, ensuring a balanced distribution, as shown in Table˜2.
@@ 4.4 User-Based Label Validation
To ensure that the attack labels in our dataset align with human perception, we conducted an IRB-approved user-based labeling task to validate our annotations. An interactive web interface was developed using Gradio [1], where each participant was shown a randomly selected raw-AR video pair from the AR-VIM dataset. Prior to starting the task, participants received a brief overview of VIM attacks in AR to calibrate their understanding. After viewing each video pair, they were asked to indicate their level of agreement with the statement: “The AR video contains a VIM attack” on a 5-point Likert scale [37], ranging from 1 (Strongly Disagree) to 5 (Strongly Agree). For non-attack cases, where disagreement indicates correct perception, we inverted the scores by subtracting them from 6, ensuring that higher values consistently indicate stronger agreement with the correct label. The user interface for data validation is shown in Fig.˜4.
We recruited 26 participants (aged 21 to 56; 8 female), with each participant evaluating 40 video pairs that were randomly selected to ensure a balanced distribution of attack labels and types. To accommodate diverse participation, the data labeling task was conducted in a hybrid format: in-person participants viewed videos on a 16-inch laptop for optimal comparison, while remote participants used the same interface displayed on their personal devices.
The distribution of agreement scores is shown in Fig.˜5. The average agreement score across the dataset is 4.53, indicating strong alignment between the dataset labels and human judgment. We also observed that visual pattern-related videos received relatively lower agreement scores, likely because humans have a more subjective and diverse understanding of visual patterns compared to text. This user-based data labeling result validates the perceptual soundness of AR-VIM and confirms that VIM attacks, as defined in our taxonomy, are generally identifiable by human users.
@@ 5 System Design
To detect VIM attacks in AR, we propose VIM-Sense, a multimodal reasoning system that operates across an edge-cloud architecture. As shown in Fig.˜6, the system takes raw and AR image pair as input, and outputs a binary prediction indicating whether an attack has occurred. VIM-Sense consists of two major components: (1) an edge server responsible for lightweight pre-processing, prompt construction, and post-processing; (2) a cloud server that performs semantic reasoning using a VLM. The edge server and the cloud server communicate through a Wi-Fi connection.
@@
After the prompt is constructed, the image encoding module encodes the image pair into Base64 format. Alongside the prompt, the encoded image pair is sent to the cloud server via an API. After the cloud server processes the task and returns the generated response, the edge server’s response post-processing module extracts the final prediction from the rich natural language explanations. We search the model’s output for the last occurrence of “Yes” or “No,” use the result as the final attack prediction, and return it to the AR interface.
@@ 6.1 Experiment Setup
Proposed Method: Our system, VIM-Sense, combines VLMs with OCR-based prompt construction and post-processing to identify attacks in raw-AR image pairs. For all experiments, we use the full pipeline described in Section˜5 unless otherwise noted. We test VIM-Sense using two state-of-the-art commercial VLMs: GPT-4o by OpenAI and Gemini-1.5-Pro by Google, as well as an open-source VLM, LLaVA-OneVision[33]. For the OCR component, we employed the EasyOCR library implemented in Python, a deep learning-based OCR engine that supports text recognition in multiple languages.
@@
OCR-only: A rule-based baseline that detects attacks by comparing the textual content extracted from the raw image () and the AR image () using the EasyOCR module. This approach does not incorporate semantic reasoning or any visual context beyond textual content. We define a text preservation threshold=0.9: if fewer than 90% of the textual elements from are preserved in , the system flags the pair as an attack. For fairness, we do not evaluate this baseline on pattern manipulation tasks, as these do not involve any textual changes.
Feature Similarity: A visual feature-based baseline that encodes both and using a pre-trained vision encoder and computes the cosine similarity between their global embeddings. A threshold=0.9 is applied to classify semantic changes: if the similarity between the embeddings is lower than the threshold, it may indicate the AR image is significantly different from the raw image and a VIM attack is likely to occur.
@@ 6.2 Results
| Feature Similarity | N/A | 50.00 | 46.77 | 62.07 | 55.00 | 47.76 | 66.10 | 46.77 | 53.32 | 0.81 | ||||||||||||
We evaluate system performance based on two key metrics: attack detection accuracy and latency.
Accuracy: Since VIM-Sense performs binary classification (attack or non-attack), we adopt standard classification accuracy as our primary metric. It is defined as the percentage of video pairs correctly identified concerning their ground truth labels. This metric directly reflects a method’s capability in detecting VIM attacks in AR.
@@
Table˜3 summarizes the detection accuracy of VIM-Sense compared with several baselines across all seven attack types, as well as the average detection latency. VIM-Sense achieves the best detection accuracy across all seven attack types, reaching an overall accuracy of 88.94% when using GPT-4o. It performs particularly well on phrase replacement (95.16%). Even in relatively low-performing categories such as pattern extra information, the system maintains a competitive performance (77.42%). These results highlight the effectiveness of combining textual grounding with semantic reasoning. Using Gemini-1.5-Pro, the overall accuracy is lower (75.00%), but still competitive given the model’s different reasoning capabilities and sensitivity to virtual content in AR. When using LLaVA-OneVision, the accuracy dropped significantly to 54.65%, suggesting that moderate-scale, resource friendly open-source VLMs are currently insufficient for effectively detecting VIM attacks.
In terms of latency, VIM-Sense requires 7.07 seconds with GPT-4o, 5.57 seconds with Gemini-1.5-Pro and 9.49 seconds with LLaVA-OneVision to return a decision after detecting virtual content. This reflects the additional processing time required for image encoding, OCR extraction, and especially multi-stage prompting. Nevertheless, it remains within an acceptable range for many AR applications, particularly in scenarios where the environment is relatively static and does not require real-time responsiveness.
@@
GenAI-Only: this baseline removes the OCR-based text grounding, relying solely on generic prompts for semantic reasoning. While it achieves a decent accuracy of 86.95% (GPT-4o), which is less than 2% lower than our proposed method, its performance on certain attack types, especially those related to text modality like phrase replacement and phrase obfuscation degrades slightly, showing that grounding the VLM in scene text improves sensitivity to text and result in a better system performance. Latency is reduced to 6.96 seconds due to simplified pipeline construction.
GenAI-Underdetailed: this baseline simplifies the prompt further by totally removing reasoning steps and directly asking the VLM if the AR content is manipulative. The lack of instruction clarity leads to a significant drop in performance ( 62.83% for GPT-4o), more than 26% lower than the proposed method. This confirms the importance of chain-of-thought guidance in understanding subtle semantic changes. It also yields a much lower latency (3.35 seconds with GPT-4o, 1.96 seconds with Gemini and 1.26 seconds with LLaVA-OneVision) because the reasoning steps are largely reduced.
OCR-Only: This baseline relies on a simple text preservation threshold to flag potential attacks. While it does not incorporate any semantic reasoning or visual pattern analysis, it surprisingly achieves a reasonable overall accuracy of 64.02%. Notably, it outperforms VIM-Sense in phrase obfuscation tasks. This is because in phrase manipulation cases, the text in the images is usually largely obstructed, while in the non-manipulated cases, the text remains largely unchanged, making this task especially well-suited for the OCR-based approach. However, it still underperforms in all other tasks compared to the proposed method, as those tasks typically involve more semantic changes without significantly altering the amount of text. For latency, OCR-Only is very fast ( 1.05 seconds) due to its simple logic and not employing a VLM.
Feature Similarity: here we encode global features from both images and compare them using cosine similarity. Despite being vision-based, it lacks semantic awareness and performs generally ordinary across all attack types, resulting in the lowest accuracy ( 53.32%). It has the lowest latency (0.81 seconds) since it only involves feature extraction and vector comparison.
@@ 6.3 Ablation Study
To better understand the role of prompt design in the performance of VIM-Sense, we conduct an ablation study focused on the system’s sensitivity to potential VIM attacks. Specifically, we examine whether adjusting the language instructions used in the prompt to reflect different assumptions about user vulnerability can influence the model’s predictions.
In the default prompt described in Section˜5, we instruct the VLM to reason under the assumption that the user is moderately resilient to deception. That is, the user is aware they are in an AR scene and will not be misled by virtual content that is stylistically inconsistent or has minimal semantic deviation from the real scene. This is reflected in the following part of the prompt template:
To explore the impact of this design choice, we evaluate two ablated versions of the prompt for attack detection. In ablation (1), we remove the sentence:
from the prompt. This eliminates the guidance about user cognitive strength, making the model more neutral. In ablation 2, we replace the entire cognitive resilience statement with:
@@
The standard prompt achieves the best overall performance, with an accuracy of 88.94%, 215 TP, and 187 TN. In Ablation 1, removing the cognitive resilience statement leads to a slight increase in FN and a significant increase in FP, resulting in a 4.4% drop in accuracy. This indicates that a minor relaxation of user modeling can reduce the system’s performance. In Ablation 2, where the prompt assumes users are easily confused, the model becomes overly sensitive to visual changes. While TP increases slightly to 228, TN drops significantly to 86, and FP jumps to 125, indicating the system is misidentifying many non-attack cases as attacks. This shift leads to the worst accuracy ( 69.47%). These findings highlight the importance of careful prompt design in the system’s sensitivity. Overly cautious prompts can trigger false alarms, while under-specified prompts may fail to capture subtle yet critical manipulations.

### v3

@@ 3.2 Data Collection Pipeline
Monitor-Based AR Data Collection: We started by collecting 58 high-resolution real-world scene images from online sources and generative AI models. The AI-generated backgrounds are a double-edged sword: they enable a broader range of scenarios that are often difficult to capture in the real world, but may also introduce domain shifts that may affect the behavior of VLMs. All of these backgrounds form a diverse set of environments such as street signs, building facades, hospital entrances, and parking instructions. We display these images on a monitor one by one as the background. For each background, we manually designed virtual content to simulate either attacked or non-attacked cases. The content varies across attack formats and purposes, based on the taxonomy introduced in Section˜2. These assets were imported into Unity and positioned at designated regions. Then, we implemented an AR application for Android phones using Unity with ARCore support. Specifically, we used ARCore’s image tracking capability to precisely align the virtual content with the background scene. In our setup, each background image is used as a reference image. When the phone’s camera detects the reference image at runtime, ARCore registers its spatial pose and enables Unity to attach virtual content to the same coordinate system. After that, we recorded each AR experience using two virtual AR cameras in Unity. One camera renders only the background scene to produce the raw image , while the other renders both the background and the virtual content to produce the AR image . Finally, we combined sequences of raw/AR images into raw/AR video pairs, each pair representing a specific scenario.
@@
Real-World AR Data Collection: While the monitor-based pipeline allows for efficient prototyping and introduces a more diverse range of scenarios, it may fail to capture important spatial properties of real-world AR environments such as 3D depth cues, which are intrinsic to real AR environments. To address this limitation, we additionally collected some data in real-world settings. Specifically, we developed an AR application for the Meta Quest 3 in Unity 2022.3.61f1, which supports access to the device’s main camera. The app allows users to grab and place virtual content at desired locations using the controller. We manually designed 35 combinations of real-world scenes and virtual content to comprehensively cover all attack formats and purposes described in Section˜2. The main camera of the Meta Quest 3 was accessed to capture raw images , while a virtual camera in Unity rendered the virtual content separately. The virtual content was then overlaid onto the raw images to produce the corresponding AR images . Finally, we also combined image sequences to create the videos.
@@ 3.3 Dataset Composition
With the data collection pipelines, we developed the AR-VIM dataset. The final AR-VIM dataset consists of 452 video pairs, each containing a raw video and its corresponding augmented version. These pairs span a total of 202 unique scenes of AR experiences. Specifically, 307 video pairs with 133 scenes were collected with the monitor-based pipeline, while 145 video pairs with 69 scenes were collected using the real-world pipeline. Each pair simulates either an attacked or non-attacked scenario using the taxonomy introduced in Section˜2. For each video pair, the virtual content appears at a specific timestamp, simulating a real-time AR interaction where the virtual overlay is introduced during scene observation. Each video in the dataset has a resolution of either 480×1080 pixels (monitor-based videos) or 960×1280 pixels (real-world videos), and a duration ranging from 4 to 17 seconds. The videos had a framerate of 15 FPS, a constraint imposed by smartphone hardware limitations, as running an AR application while simultaneously capturing high-frame-rate raw and AR videos places strain on AR devices.
To ensure balanced coverage across both manipulation formats and purposes, we constructed seven types of attacks, as elaborated upon in Section˜2.3: character replacement, phrase replacement, phrase obfuscation, phrase extra information, pattern replacement, pattern obfuscation, and pattern extra information. For each attack type, we also generated a comparable number of non-attacked cases for comparison. In total, the dataset includes 241 attacked video pairs and 211 non-attacked video pairs, ensuring a balanced distribution, as shown in Table˜2.
@@ 3.4 User-Based Label Validation
To ensure that the attack labels in our dataset align with human perception, we conducted an IRB-approved user-based labeling task to validate our annotations. An interactive web interface was developed using Gradio [1], where each participant was shown a randomly selected raw-AR video pair from the AR-VIM dataset. Prior to starting the task, participants received a brief overview of VIM attacks in AR to calibrate their understanding. After viewing each video pair, they were asked to indicate their level of agreement with the statement: "The AR video contains a VIM attack" on a 5-point Likert scale [38], ranging from 1 (Strongly Disagree) to 5 (Strongly Agree). For non-attacked cases, where disagreement indicates correct perception, we inverted the scores by subtracting them from 6, ensuring that higher values consistently indicate stronger agreement with the correct label. The user interface for data validation is shown in Fig.˜4.
We recruited 26 participants (aged 21 to 56; 8 female), with each participant evaluating 40 video pairs that were randomly selected to ensure a balanced distribution of attack labels and types. To support both in-person and remote participation, the data labeling task was conducted in a hybrid format: in-person participants viewed videos on a 16-inch laptop, while remote participants used the same interface displayed on their personal devices.
The distribution of agreement scores is shown in Fig.˜5. The average agreement score across the dataset is 4.53, indicating strong alignment between the dataset labels and human judgment. We also observed that visual pattern-related videos received relatively lower agreement scores, likely because humans have a more subjective and diverse understanding of visual patterns compared to text. The user-based data labeling results validate the perceptual soundness of AR-VIM and confirms that VIM attacks, as defined in our taxonomy, are generally identifiable by human users.
@@ 4 System Design
To detect VIM attacks in AR, we propose VIM-Sense, a multimodal reasoning system that operates across an edge-cloud architecture. As shown in Fig.˜6, the system takes a raw and AR image pair as input, and outputs a binary prediction indicating whether an attack has occurred. VIM-Sense consists of two major components: (1) an edge server responsible for lightweight pre-processing, prompt construction, and post-processing; (2) a cloud server that performs semantic reasoning using a VLM. The edge server accesses the models on the cloud over the internet via Wi-Fi.
@@
After the prompt is constructed, the image encoding module encodes the image pair into Base64 format. Alongside the prompt, the encoded image pair is sent to the cloud server via an API. After the cloud server processes the task and returns the generated response, the edge server’s response post-processing module extracts the final prediction from the rich natural language explanations. We search the model’s output for the last occurrence of "Yes" or "No," use the result as the final attack prediction, and return it to the AR interface.
@@ 5.1 Experiment Setup
Proposed Method: Our system, VIM-Sense, combines VLMs with OCR-based prompt construction and post-processing to identify attacks in raw-AR image pairs. For all experiments, we use the full pipeline described in Section˜4 unless otherwise noted. We test VIM-Sense using two state-of-the-art commercial VLMs, GPT-4o by OpenAI and Gemini-1.5-Pro by Google, as well as an open-source VLM, LLaVA-OneVision[34]. For the OCR component, we employed the EasyOCR library implemented in Python, a deep learning-based OCR engine that supports text recognition in multiple languages.
@@
OCR-only: A rule-based baseline that detects attacks by comparing the textual content extracted from the raw image and the AR image using the EasyOCR module. This approach does not incorporate semantic reasoning or any visual context beyond textual content. We define a text preservation threshold of 0.9: if fewer than 90% of the textual elements from are preserved in , the system flags the pair as an attack. For fairness, we do not evaluate this baseline on pattern manipulation tasks, as these do not involve textual changes.
Feature Similarity: A visual feature-based baseline that encodes both and using a pre-trained vision encoder and computes the cosine similarity between their global embeddings. A threshold of 0.9 is applied to classify semantic changes: if the similarity between the embeddings is lower than the threshold, it may indicate the AR image is significantly different from the raw image, and a VIM attack is likely to occur.
@@ 5.2 Results
Feature Similarity | N/A | 50.00 | 46.77 | 62.07 | 55.00 | 47.76 | 66.10 | 46.77 | 53.32 | 0.81 |
We evaluate system performance using two key metrics: attack detection accuracy and latency.
Accuracy: Since VIM-Sense performs binary classification (attack or non-attack), we adopt classification accuracy as our primary metric. It is defined as the percentage of video pairs correctly identified with respect to their ground truth labels. This metric directly reflects a method’s capability in detecting VIM attacks in AR.
@@
Table˜3 summarizes the detection accuracy of VIM-Sense compared with several baselines across all seven attack types, as well as the average detection latency. VIM-Sense achieves the best detection accuracy across all seven attack types, reaching an overall accuracy of 88.94% when using GPT-4o. It performs particularly well on phrase replacement (95.16%). Even in relatively low-performing categories such as pattern extra information, the system maintains a competitive performance (77.42%). These results highlight the effectiveness of combining textual grounding with semantic reasoning. Using Gemini-1.5-Pro, the overall accuracy is lower (75.00%), but still competitive given the model’s different reasoning capabilities and sensitivity to virtual content in AR. When using LLaVA-OneVision, the accuracy dropped significantly to 54.65%, suggesting that moderate-scale, resource-friendly open-source VLMs are currently insufficient for effectively detecting VIM attacks.
In terms of latency, VIM-Sense requires 7.07 seconds with GPT-4o, 5.57 seconds with Gemini-1.5-Pro, and 9.49 seconds with LLaVA-OneVision to return a decision after detecting virtual content. This reflects the additional processing time required for image encoding, OCR extraction, and especially multi-stage prompting. Nevertheless, it remains within an acceptable range for many AR applications, particularly in scenarios where the environment is relatively static and does not require real-time responsiveness.
@@
GenAI-Only: This baseline removes the OCR-based text grounding, relying solely on generic prompts for semantic reasoning. While it also achieves a high accuracy of 86.95% (GPT-4o), which is less than 2% lower than our proposed method, its performance on certain attack types, especially those related to text modality like phrase replacement and phrase obfuscation degrades slightly, showing that grounding the VLM in scene text improves sensitivity to text and results in a better system performance. Latency is reduced to 6.96 seconds due to simplified pipeline construction.
GenAI-Underdetailed: This baseline simplifies the prompt further by completely removing reasoning steps and directly asking the VLM if the AR content is manipulative. The lack of instruction clarity leads to a significant drop in performance ( 62.83% for GPT-4o, more than 26% lower than the proposed method.) This confirms the importance of chain-of-thought guidance in understanding subtle semantic changes. It also yields a much lower latency (3.35 seconds with GPT-4o, 1.96 seconds with Gemini and 1.26 seconds with LLaVA-OneVision) because the reasoning steps are largely reduced.
OCR-Only: This baseline relies on a simple text preservation threshold to flag potential attacks. While it does not incorporate any semantic reasoning or visual pattern analysis, it surprisingly achieves a moderate overall accuracy of 64.02%, which outperforms all the GenAI-Underdetailed baselines. Notably, it outperforms VIM-Sense in phrase obfuscation tasks. This is because in phrase obfuscation cases, the text in the images is usually largely obstructed, while in the non-obfuscated cases, the text remains largely unchanged, making this task especially well-suited for the OCR-based approach. However, it still underperforms in all other tasks compared to the proposed method, as those tasks typically involve more semantic changes without significantly altering the length of text. With a latency of 1.05 seconds, OCR-Only is very fast due to its simple logic and not employing a VLM.
Feature Similarity: Here we encode global features from both images and compare them using cosine similarity. Despite being vision-based, this baseline lacks semantic awareness and performs weakly across all attack types, resulting in the lowest accuracy ( 53.32%). It has the lowest latency (0.81 seconds) since it only involves feature extraction and vector comparison.
@@ 5.3 Ablation Study
To better understand the role of prompt design in the performance of VIM-Sense, we conduct an ablation study focusing on the system’s sensitivity to potential VIM attacks. Specifically, we examine whether adjusting the language instructions used in the prompt to reflect different assumptions about user vulnerability can influence the model’s predictions.
In the default prompt described in Section˜4, we instruct the VLM to reason under the assumption that the user is moderately resilient to deception. That is, the user is aware they are in an AR scene and will not be misled by virtual content that is stylistically inconsistent or has minimal semantic deviation from the real scene. This is reflected in the following part of the prompt template:
To explore the impact of this design choice, we evaluate two ablated versions of the prompt for attack detection. In ablation 1, we remove the sentence:
from the prompt. This eliminates the guidance about the cognitive strength of the user, making the model more neutral. In ablation 2, we replace the entire cognitive resilience statement with:
@@
The standard prompt achieves the best overall performance, with an accuracy of 88.94%, 215 TP, and 187 TN. In ablation 1, removing the cognitive resilience statement leads to a slight increase in FN and a significant increase in FP, resulting in a 4.4% drop in accuracy. This indicates that a minor relaxation of user modeling can reduce the system’s performance. In ablation 2, where the prompt instructs the VLM to assume that users are easily confused, the model becomes overly sensitive to visual changes. While TP increases slightly to 228, TN drops significantly to 86, and FP jumps to 125, indicating that the system is misidentifying many non-attacked cases as attacks. This shift leads to the worst accuracy ( 69.47%). These findings highlight the importance of careful prompt design in affecting the system’s sensitivity to potential attacks. Overly cautious prompts can trigger false alarms, while under-specified prompts may fail to capture subtle yet critical manipulations.